OpenStudy (anonymous):

Find the slope of the curve f(x)=(x^2)-3 at P(2,1)

OpenStudy (wolf1728):

We solve this problem by taking the derivative of x² -3 which is 2x at point P, x = 2 and so the slope is 2 • 2 or 4

OpenStudy (anonymous):

\[f(x)=x^2-3\]\[f'(x)=2x=m\] At the point (2,1): \[f'(2)=2(2)=4=m\]Therefore the slope of f(x) at (2,1) is 4.